OpenStudy (anonymous):

Solve the system by using elimination or elimination with multiplication. x + y = 16 and x + y = -2

OpenStudy (mathstudent55):

To use elimination, you need to add the equations and make one variable add up to zero. Since you have x in both equations, multiply the second equation by -1 to get -x. Then when you add the equations, the x's will be eliminated becasue x + (-x ) = 0 x + y = 16 x + y = -2 Rewrite first equation. Below that write second equation multiplied by -1: x + y = 16 -x - y = 2 (add equations) --------------- 0 + 0 = 18 0 = 18 Since 0 is not equal to 18, this system of equations has no solution.

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (mathstudent55):

You're welcome.